OpenStudy (anonymous):

@jim_thompson5910 I'm stuck between B and D ! :(

jimthompson5910 (jim_thompson5910):

this sort of thing involves a lot of drawings and it can get very messy hopefully this page makes things simple (or at least less complicated) http://www.math.nmsu.edu/~breakingaway/Lessons/PTUST/PTUST.html let me know if that page helps or not

OpenStudy (anonymous):

Confused..

jimthompson5910 (jim_thompson5910):

where at?

OpenStudy (anonymous):

divisions on bottom

jimthompson5910 (jim_thompson5910):

you mean like the c in a/c ?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

do you see how c is the hypotenuse of the largest triangle?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

also notice how they isolated triangle 1 and triangle 2

jimthompson5910 (jim_thompson5910):

they drew them below the largest triangle

jimthompson5910 (jim_thompson5910):

for triangle 1, a is the hypotenuse for the largest triangle, c is the hypotenuse they correspond which means a/c ------------------------------------------------------- at the same time, for triangle 1, x is the short leg for the largest triangle, a is the short leg they correspond ---> x/a ------------------------------------------------------- the two ratios are equal which is how they got a/c = x/a

jimthompson5910 (jim_thompson5910):

make sense?

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

they use a similar approach to get b/c = y/b make sure you understand how they got that

jimthompson5910 (jim_thompson5910):

be sure to focus on triangle 2 and the largest triangle

OpenStudy (anonymous):

>.<

jimthompson5910 (jim_thompson5910):

still stuck?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

focus on triangle 2 and the largest triangle for triangle 2, what is b? hypotenuse? short leg? long leg?

OpenStudy (anonymous):

b is the long leg I went back and re-read your response. Yes, it makes sense.

jimthompson5910 (jim_thompson5910):

b isn't the long leg for triangle 2

jimthompson5910 (jim_thompson5910):

notice how it's opposite the 90 degree angle

OpenStudy (anonymous):

where is the 90 degree angle?

jimthompson5910 (jim_thompson5910):

it's where segment h and segment y meet

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

yes that is correct

jimthompson5910 (jim_thompson5910):

those angles you marked are all 90 degrees

jimthompson5910 (jim_thompson5910):

for triangle 2, b is opposite the 90 degree angle, so b is actually the hypotenuse longest side is always opposite largest angle

OpenStudy (anonymous):

Yahooo! :) It makes a bit more sense with the 90 degree angles marked.

jimthompson5910 (jim_thompson5910):

sorry I had them flipped, but here it's fixed

jimthompson5910 (jim_thompson5910):

based off that, we see that triangle ADB ~ triangle CDA (look at triangle 1 & 2, notice how they match up)

jimthompson5910 (jim_thompson5910):

we also see that triangle BCA ~ triangle BAD (focus on larger triangle and triangle 1)

OpenStudy (anonymous):

Okay...

jimthompson5910 (jim_thompson5910):

so were you able to get any further based on that?

OpenStudy (anonymous):

I think it's B.

jimthompson5910 (jim_thompson5910):

why is that?

OpenStudy (anonymous):

A =D

jimthompson5910 (jim_thompson5910):

I don't know what you mean

jimthompson5910 (jim_thompson5910):

how are A and D equal?

OpenStudy (anonymous):

A=D oh crap but C is also equivalent to A.. -_-

jimthompson5910 (jim_thompson5910):

do you see how triangle ADB ~ triangle CDA and triangle BCA ~ triangle BAD ??

OpenStudy (anonymous):

yeah.

jimthompson5910 (jim_thompson5910):

I'm going to call the overall largest triangle, triangle 3 triangles 1 and 2 are below triangle 3

jimthompson5910 (jim_thompson5910):

triangle ADB and triangle CDA are what two triangles? (use labels triangle 1, triangle 2 or triangle 3)

jimthompson5910 (jim_thompson5910):

triangle ADB is _______ fill in the blank with triangle 1, triangle 2 or triangle 3 (look at how I labeled triangle ADB)

OpenStudy (anonymous):

1

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

DAC is 2

jimthompson5910 (jim_thompson5910):

triangle CDA is triangle 2, yes

jimthompson5910 (jim_thompson5910):

think back to this page http://www.math.nmsu.edu/~breakingaway/Lessons/PTUST/PTUST.html notice how they connect triangle 1 with triangle 3 (to get a/c = x/a) and triangle 2 with triangle 3 (to get b/c = y/b)

jimthompson5910 (jim_thompson5910):

they do NOT connect triangle 1 with triangle 2

jimthompson5910 (jim_thompson5910):

so triangle ADB ~ triangle CDA is definitely true, but it's not relevant or used here

jimthompson5910 (jim_thompson5910):

that allows us to eliminate the first two answer choices

jimthompson5910 (jim_thompson5910):

make sense?

OpenStudy (anonymous):

Y E S : D

jimthompson5910 (jim_thompson5910):

let's look at choice D does BA correspond to DA (look at triangle 3 and triangle 1) ?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

BA corresponds to BD instead, so choice D is also eliminated

OpenStudy (anonymous):

Ah, it is C.

OpenStudy (anonymous):

:D !!

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Thank you!!

jimthompson5910 (jim_thompson5910):

np